# DESY-25-105

### v1

@@ 3.1 $(A_{1},A_{2})$ theory
Now, assign bi-degree to $x^{(\alpha)}$ as $\textrm{deg}(x^{(\alpha)})=(2,\alpha)$; then, one can see that $\textrm{deg}(f^{(\alpha)})=(4,\alpha)$. The arc space of the double point is given as
@@
To compute the Hilbert series, notice that the arc space of the polynomial ring over $X$ can be decomposed according to this bi-degree to get
@@
Now, at each order in $n$, the number of generators is identical to the number of partitions of $n$ into elements $1$ and $2$. The generating function of such partitions is given as $1/(1-q)(1-q^{2})$. At order $q^{n}p^{4}$, we also notice that one relation, $f^{(n)}$, kills one of the generators, so that we get the contribution of $-q^{n}p^{4}$. Summing up in $n$, we get
@@ 3.2 $(A_{1},A_{4})$ theory
and where the degree of the single coordinate $x$ is $\operatorname{deg}(x)=2$. Again one can write the decomposition of the arc space according to the bi-grading as we did for $(A_{1},A_{2})$ in equation (3.3). Then, we simply need to compute the dimensions of the vector spaces $V_{n,m}$ taking into account the relations and syzygys. Instead of proceeding by hand, we can ask Macaulay2 to determine a Gröbner basis and perform the calculation for jet schemes of sufficiently low jet degree. As discussed already, we know that
@@
$\operatorname{HS}_{q,q}(J_{\infty}(R))=\sum_{i=0}^{14}a_{i}q^{i}+\mathcal{O}(q^{15})\,,$ (3.15)
@@ 3.3 $(A_{1},A_{6})$ theory
[…]
@@
$(1,0,1,1,2,2,4,4,6,7,10,12,17,19,26)\,,$ (3.20)
as determined via truncation to the $10$th jet scheme. In these $(A_{1},A_{2n})$ examples, it is straightforward to go to much higher orders numerically using Macaulay2; while we have verified that these are consistent with our proposal, we do not write the full expressions here for brevity.
@@ 3.4 $(A_{2},A_{3})$ theory
$X=\operatorname{Spec}(\mathbb{C}[x,y]/(x^{2}-y^{3},xy))\,,$ (3.21)
@@
where $I_{n}$ is the ideal generated by the $22n+$ 2 polynomials,
$\sum_{\begin{subarray}{c}i,j\geq 0\\[1.39998pt] i+j=m\end{subarray}}x_{i}x_{j}-\sum_{\begin{subarray}{c}i,j,k\geq 0\\[1.39998pt] i+j+k=m\end{subarray}}y_{i}y_{j}y_{k}\ ,\qquad\sum_{\begin{subarray}{c}i,j\geq 0\\[1.39998pt] i+j=m\end{subarray}}x_{i}y_{j}$ $\left(m=0,\cdots,n\right).$ (3.24)
Recalling that $x_{i}$ has bi-degree $(3,i)$ and $y_{i}$ has bi-degree $(2,i)$, we can now determine the bi-graded Hilbert series for $J_{n}(R)$ by brute-force. We find that
@@ 3.5 $(A_{2},A_{4})$ theory
$X=\operatorname{Spec}(\mathbb{C}[x,y]/(x^{2}-y^{3},xy^{2}))\,,$ (3.27)
@@ 3.6 $(A_{2},A_{6})$ theory
The analysis for the $(A_{2},A_{6})$ theory follows in largely the same vein as the $(A_{2},A_{4})$ theory. Our affine scheme was determined in Section 2, and it is
$X=\operatorname{Spec}(\mathbb{C}[x,y]/(x(x^{2}-y^{3}),y(x^{2}-y^{3}))\,.$ (3.31)
The variables $x$ and $y$ are also graded in the same way as for $(A_{2},A_{4})$ as given in equation (3.28). Now we can write down the coordinates and ideal defining the arc space, and we can compute the Schur limit of the Hilbert series to a particular order by truncating to the appropriate jet scheme. We find that
$\operatorname{HS}(J_{\infty}(R))(q,q)=\sum_{j=0}^{15}a_{j}q^{j}+\mathcal{O}(q^{16})\,,$ (3.32)
@@
$(1,0,1,2,3,4,8,10,16,22,33,45,66,86,122,163)\,.$ (3.33)
The coefficients are again determined via an explicit computation of the Hilbert series of the $13$th jet scheme using Macaulay2. These again match with the expansion of the Schur index for the $(A_{2},A_{6})$ theory as given in equation (2.19).
@@ 3.7 $(A_{2},A_{7})$ theory
In the final explicit example of the $(A_{2},A_{N-1})$ sequence of Argyres–Douglas SCFTs, we consider the $(A_{2},A_{7})$ theory. Our affine scheme was determined from the associated vertex operator algebra in Section 2; it is
$X=\operatorname{Spec}(\mathbb{C}[x,y]/(x(x^{2}-y^{3}),y^{2}(x^{2}-y^{3}))\,,$ (3.34)
where the variables $x$ and $y$ have the same grading as given in equation (3.28). Considering the $13$th jet scheme, we can again determine the leading orders of the Schur limit of the Hilbert series of the arc space using Macaulay2. We find that the coefficients are
$(1,0,1,2,3,4,8,10,17,23,34,47,70,92,132,175)\,,$ (3.35)
@@ 3.8 $(A_{3},A_{4})$ theory
We now consider the example of the $(A_{3},A_{4})$ Argyres–Douglas theory. Our affine scheme, as worked out in Section 2, is
$X=\operatorname{Spec}(\mathbb{C}[x,y,z]/(xz+y^{2}+z^{3},xy+yz^{2},x^{2}+y^{2}z+z^{4})\,,$ (3.36)
@@
Based on the degree of $z$ being $2$, we can use the jet scheme $J_{9}(R)$ to determine the coefficients of the Schur limit of the Hilbert series of the arc space up to and including order $q^{11}$. Again utilizing Macaulay2, we find that the leading coefficients are
$(1,0,1,2,4,5,9,12,21,29,44,59)\,.$ (3.38)
@@ 4 Macdonald index
Effectively, this way of introducing an extra grading is tantamount to counting the number of letters needed to construct a (basis) vector in $\textrm{gr}V$ endowed with grading $w$. When there is more than one type of letter, then the relation may not respect this grading based on the degree of letters. For example, when we have $x^{2}=y^{3}$ with $w(x)=2$ and $w(x)=1$, the LHS has $w=4$ whereas RHS has $w=3$. The relation implies that we should not count both.
Our prescription is simply that whenever the relation mixes the $T$-grading, drop the states with the highest $T$-weight. This is realized by equation (4.3). For example, suppose the relation identifies the states with $T$-weight of $m$ and $m-1$. Then we see that such a vector of $T$-weight $m$ should be absent in $\textrm{gr}V^{(m)}=V_{m}/V_{m-1}$ from the relation. However such a vector is not removed in $\textrm{gr}V^{(m-1)}=V_{m-1}/V_{m-2}$ since the relation is in $V_{m}$ but not in $V_{m-1}$.
@@ 5.1 $\mathbb{C}^{2}/\mathbb{Z}_{2}$
Therefore, we add an additional coordinate that plays the role of the stress-tensor. To this end, we start with the following coordinate ring:
$R=\mathbb{C}[x,y,z,w]/(w-x^{2}-y^{2}-z^{2})\,.$ (5.3)
Note that this does not reduce to the $\mathbb{C}^{2}/\mathbb{Z}_{2}$ singularity as a variety. This sets the coordinates to have degrees
Upon computing the Hilbert series of the arc space (using Macaulay2), we find
$\operatorname{HS}_{q,q}(J_{\infty}(X))=1+3q+9q^{2}+22q^{3}+51q^{4}+108q^{5}+221q^{6}+429q^{7}+810q^{8}+\cdots\ .$ (5.5)
This expression does not appear to agree with any known 4d SCFT. Instead, we observe that this Hilbert series agrees with the Schur index of the $SU(2)$ current multiplet, given by
$\operatorname{PE}\left[\frac{q}{1-q}\chi_{3}\right]=1+q\chi_{3}+q^{2}(\chi_{1}+\chi_{3}+\chi_{5})+q^{3}(\chi_{1}+3\chi_{3}+\chi_{5}+\chi_{7})+\cdots\,,$ (5.6)
where $\chi_{n}$ denotes the character for the $n$-dimensional representation of the $SU(2)$ flavor symmetry. [^16] This index does not agree with that of the free hypermultiplet with $\mathbb{Z}_{2}$ flavor symmetry gauged, which is given as $I=1+3q+9q^{2}+22q^{3}+46q^{4}+93q^{5}+176q^{6}+\cdots$. It would be interesting to find an affine scheme that reproduces this simple theory.
@@
We can see that once $X$ is reduced, we obtain the $\mathbb{C}^{2}/\mathbb{Z}_{2}$ singularity as a variety. We can use Macaulay2 to determine the Schur limit of the Hilbert series of the $4$th jet scheme, which should overlap with the Schur limit of the Hilbert series of the arc space up to and including the $q^{5}$ term. We find:
$\operatorname{HS}_{q,q}(J_{\infty}(R))=1+3q+9q^{2}+19q^{3}+42q^{4}+81q^{5}+\cdots\,.$ (5.9)
@@
$\operatorname{PE}\left[\frac{q-q^{3}}{(1-q)(1-q^{3})}\chi_{3}\right]=1+3q+9q^{2}+19q^{3}+42q^{4}+81q^{5}+155q^{6}+276q^{7}+\cdots\,,$ (5.10)
where we have turned off the flavor fugacity for simplicity. We see that these match perfectly at least up to and including order $q^{5}$. While it would be useful to go to higher orders to provide a more convincing demonstration of our conjecture in equation (1.4), Macaulay2 has difficulty finding Gröbner bases for the higher jet schemes. In the recent work [5], the authors proved that the arc space of our affine scheme indeed gives the Schur and Macdonald indices.
Let us make a comment about another possible choice for the affine scheme. Consider the following scheme:
$X=\operatorname{Spec}\left(\mathbb{C}[x,y,z]/(x(x^{2}+y^{2}+z^{2}),y(z^{2}+y^{2}+z^{2}),z(x^{2}+y^{2}+z^{2}))\right)\,,$ (5.11)
which again reduces to the same $\mathbb{C}^{2}/\mathbb{Z}_{2}$ singularity. However, the Hilbert series of the arc space is now given by
$\operatorname{HS}_{q,q}(J_{\infty}(R))=1+3q+9q^{2}+19q^{3}+42q^{4}+79q^{5}+\cdots\,,$ (5.12)
which is not identical to the Schur index of the $(A_{1},A_{3})=(A_{1},D_{3})=\mathcal{D}_{3}(SU(2))$ theory. Thus, we see the importance of the additional coordinate $w$, which implements the Sugawara relation between stress-tensor and the currents at the level of VOA. [^17] We note that Zhu’s $C_{2}$-algebra for the $(A_{1},D_{3})$ theory, as determined in [17], is the polynomial ring of the scheme $X$ in equation (5.11): $R_{V}=\mathbb{C}[X]$. Thus, it appears that our affine scheme, which captures the Schur and Macdonald indices, is not generically identical to the spectrum of Zhu’s $C_{2}$-algebra of the associated VOA. Recently, [5] has proposed an alternative Zhu’s $C_{2}$-algebra for the $(A_{1},D_{3})$ theory, which incorporates the stress-energy tensor, and whose spectrum thus gives the same affine scheme as we proposed in equation (5.7). It would be interesting to understand this apparent discrepancy.
@@
$\operatorname{HS}_{q,q}(J_{\infty}(R))=1+3q+9q^{2}+22q^{3}+51q^{4}+105q^{5}+212q^{6}+402q^{7}+744q^{8}+\cdots\,,$ (5.14)
@@
up to and including the $q^{8}$ term. Therefore, we conjecture that the following affine scheme captures the Macdonald and Schur indices of the $(A_{1},D_{2n+1})=\mathcal{D}_{2n+1}(SU(2))$ theory:
@@ 5.2 $\mathbb{C}^{2}/\mathbb{Z}_{n+1}$
which is known to have $\mathbb{C}^{2}/\mathbb{Z}_{3}$ as its Higgs branch. The associated vertex operator algebra for this theory is given by the Bershadsky–Polyakov (BP) W-algebra [21, 69], whose connection to Argyres–Douglas theories was found in [39]. The BP algebra has strong generators of dimensions $1,3/2,3/2,2$, and its Zhu’s algebra $R_{V}$ is identical to our $R$ so that the associated scheme $X_{V}$ is identical to our affine scheme $X$ [^18] There is a conjectural form of the Zhu’s algebra written in [80] for the $(A_{1},A_{2n+1})$ theory, which is slightly different from what we have. This gives a different Hilbert series which does not match with the Schur index.
Next, we can consider an affine scheme which reduces to the variety $\mathbb{C}^{2}/\mathbb{Z}_{4}$. A reasonably obvious candidate, given equation (5.18) for $\mathbb{C}^{2}/\mathbb{Z}_{3}$, is
$\mathbb{C}[x,y,z,t]/\left(xy+z^{4}-tz^{2},xt,yt,z(xy+z^{4})\right)\,,$ (5.21)
where the degrees of $(x,y,z,t)$ are $(2,2,1,2)$. The Schur limit of the Hilbert series of the arc space of $R=\mathbb{C}[X]$ can be computed to give
$\operatorname{HS}_{q,q}(J_{\infty}(R))=1+q+5q^{2}+9q^{3}+20q^{4}+36q^{5}+71q^{6}+119q^{7}+\cdots\,.$ (5.22)
This result agrees with the Schur index of $(A_{1},A_{7})$ theory [76]
$I_{(A_{1},A_{7})}=\operatorname{PE}\left[\frac{q+q^{2}\left(z^{2}+\frac{1}{z^{2}}+1\right)-q^{4}\left(z^{2}+\frac{1}{z^{2}}+1\right)-q^{5}}{(1-q)(1-q^{5})}\right]\ ,$ (5.23)
to leading orders in $q$, and where the flavor fugacity $z$ is set to one. As before, we find our affine scheme $X$ to be identical to the associated scheme $X_{V}$ of the generalized Bershadsky–Polyakov algebra [17].
Given this analysis, it seems natural to conjecture that the affine scheme that we want to associate to the $(A_{1},A_{2n-1})$ Argyres–Douglas theory is the spectrum of
$R=\mathbb{C}[x,y,z,t]/\left(xy+z^{n}-tz^{n-2},xt,yt,z(xy+z^{n})\right)\,,$ (5.24)
where the degrees of the coordinates are
$\operatorname{deg}(x)=\operatorname{deg}(y)=\frac{n}{2}\,,\quad\operatorname{deg}(z)=1\,,\quad\operatorname{deg}(t)=2\,.$ (5.25)
For larger values of $n$, the lower orders of the Hilbert series of the arc space actually become easier to compute explicitly, as the relations in the ideal get pushed to higher-orders, and thus become irrelevant for the lower-order terms in the series. As a simple, higher-rank example, the Schur limit of the Hilbert series of the ring $R$ with $n=8$ is
$\operatorname{HS}_{q,q}(J_{\infty}(R))=1+q+3q^{2}$ $+5q^{3}+12q^{4}+20q^{5}+37q^{6}$ (5.26)
$+61q^{7}+107q^{8}+171q^{9}+281q^{10}+\cdots\,,$
which matches with the Schur index of the $(A_{1},A_{15})$ theory, to the computed/written order.

### v2

@@ 3.1 $(A_{1},A_{2})$ theory
Now, assign bidegree to $x^{(\alpha)}$ as $\operatorname{deg}(x^{(\alpha)})=(2,\alpha)$; then, one can see that $\operatorname{deg}(f^{(\alpha)})=(4,\alpha)$. The arc space of the double point is given as
@@
To compute the Hilbert series, notice that the arc space of the polynomial ring over $X$ can be decomposed according to this bidegree to get
@@
Now, at each order in $n$, the number of generators is identical to the number of partitions of $n$ into elements $1$ and $2$. The generating function of such partitions is given as $1/(1-q)(1-q^{2})$. At order $q^{n}p^{4}$, we also notice that one relation, $f^{(n)}$, kills one of the generators, so that we get the contribution of $q^{n}p^{4}$. Summing up in $n$, we get
@@ 3.2 $(A_{1},A_{4})$ theory
and where the degree of the single coordinate $x$ is $\operatorname{deg}(x)=2$. Again one can write the decomposition of the arc space according to the bigrading as we did for $(A_{1},A_{2})$ in equation (3.3). Then, we simply need to compute the dimensions of the vector spaces $V_{n,m}$ taking into account the relations and syzygys. Instead of proceeding by hand, we can ask Macaulay2 to determine a Gröbner basis and perform the calculation for jet schemes of sufficiently low jet degree. As discussed already, we know that
@@
$\operatorname{HS}_{q,q}(J_{\infty}(R))=\sum_{i=0}^{16}a_{i}q^{i}+\mathcal{O}(q^{17})\,$, (3.15)
@@ 3.3 $(A_{1},A_{6})$ theory
$\begin{aligned} \operatorname{HS}_{p,q}(J_{\infty}(R))&=1+p^{2}+p^{4}+p^{6}+\left(p^{2}+p^{4}+p^{6}\right)q+\left(p^{2}+2p^{4}+2p^{6}+p^{8}\right)q^{2}\\ &\ +\left(p^{2}+2p^{4}+3p^{6}+2p^{8}\right)q^{3}+\left(p^{2}+3p^{4}+4p^{6}+4p^{8}+p^{10}\right)q^{4}\\ &\ +\left(p^{2}+3p^{4}+5p^{6}+5p^{8}+2p^{10}\right)q^{5}+\left(p^{2}+4p^{4}+7p^{6}+8p^{8}+4p^{10}+p^{12}\right)q^{6}\\ &\ +\left(p^{2}+4p^{4}+8p^{6}+10p^{8}+5p^{10}+p^{12}\right)q^{7}+\left(p^{2}+5p^{4}+10p^{6}+14p^{8}\right.\\ &\quad\;\left.+10p^{10}+2p^{12}\right)q^{8}+\cdots\,.\end{aligned}$
@@
(1,0,1,1,2,2,4,4,6,7,10,12,17,19,26,31,40,47),(1,0,1,1,2,2,4,4,6,7,10,12,17,19,26,31,40,47)\,, (3.20)
as determined via truncation to the $15$th jet scheme. In these $(A_{1},A_{2n})$ examples, it is straightforward to go to much higher orders numerically using Macaulay2; while we have verified that these are consistent with our proposal, we do not write the full expressions here for brevity.
@@ 3.4 $(A_{2},A_{3})$ theory
$X=\operatorname{Spec}(\mathbb{C}[x,y]/(x^{2}+y^{3},xy))\,$, (3.21)
@@
where $I_{n}$ is the ideal generated by the $2n+2$ polynomials,
$\sum_{\begin{subarray}{c}i,j\geq 0\\[1.63885pt] i+j=m\end{subarray}}x_{i}x_{j}-\sum_{\begin{subarray}{c}i,j,k\geq 0\\[1.63885pt] i+j+k=m\end{subarray}}y_{i}y_{j}y_{k}\ ,\qquad\sum_{\begin{subarray}{c}i,j\geq 0\\[1.63885pt] i+j=m\end{subarray}}x_{i}y_{j}$ $\left(m=0,\cdots,n\right)\,.$ (3.24)
Recalling that $x_{i}$ has bidegree $(3,i)$ and $y_{i}$ has bidegree $(2,i)$, we can now determine the bigraded Hilbert series for $J_{n}(R)$ by brute-force. We find that
@@ 3.5 $(A_{2},A_{4})$ theory
$X=\operatorname{Spec}(\mathbb{C}[x,y]/(x^{2}+y^{3},xy^{2}))\,$, (3.27)
@@ 3.6 $(A_{2},A_{6})$ theory
The analysis for the $(A_{2},A_{6})$ theory follows in largely the same vein as the $(A_{2},A_{4})$ theory. Our affine scheme was mentioned in Section 2; the degrees of the polynomials were fixed by the known orders of the null relations in the associated W-algebra, and the coefficients are fixed by requiring the reproduction of the Schur index when computing the Hilbert series of the arc space. In this way, we find the affine scheme
$X=\operatorname{Spec}(\mathbb{C}[x,y]/(x(x^{2}+4y^{3}),y(3x^{2}+y^{3}))\,.$ (3.31)
The variables $x$ and $y$ are graded in the same way as for $(A_{2},A_{4})$ as given in equation (3.28). Now we can write down the coordinates and ideal defining the arc space, and we can compute the Schur limit of the Hilbert series to a particular order by truncating to the appropriate jet scheme. We find that
$\operatorname{HS}(J_{\infty}(R))(q,q)=\sum_{j=0}^{16}a_{j}q^{j}+\mathcal{O}(q^{17})\,$, (3.32)
@@
(1,0,1,2,3,4,8,10,16,22,33,44,66,86,121,162,221).(1,0,1,2,3,4,8,10,16,22,33,44,66,86,121,162,221)\,. (3.33)
The coefficients are again determined via an explicit computation of the Hilbert series of the $14$th jet scheme using Macaulay2. These again match with the expansion of the Schur index for the $(A_{2},A_{6})$ theory as given in equation (2.20).
@@ 3.7 $(A_{2},A_{7})$ theory
In the final explicit example of the $(A_{2},A_{N-1})$ sequence of Argyres–Douglas SCFTs, we consider the $(A_{2},A_{7})$ theory. Similarly to the $(A_{2},A_{6})$ case, the polynomials appearing in our affine scheme were determined, up to coefficients, from the associated vertex operator algebra in Section 2, and the coefficients are fixed via the matching with the Schur index. The affine scheme is
$X=\operatorname{Spec}(\mathbb{C}[x,y]/(x(2x^{2}+3y^{3}),y^{2}(5x^{2}+y^{3}))\,$, (3.34)
where the variables $x$ and $y$ have the same grading as given in equation (3.28). Considering the $15$th jet scheme, we can again determine the leading orders of the Schur limit of the Hilbert series of the arc space using Macaulay2. We find that the coefficients are
(1,0,1,2,3,4,8,10,17,23,34,47,70,92,132,177,243,322),(1,0,1,2,3,4,8,10,17,23,34,47,70,92,132,177,243,322)\,, (3.35)
@@ 3.8 $(A_{3},A_{4})$ theory
We now consider the example of the $(A_{3},A_{4})$ Argyres–Douglas theory. Our affine scheme, as worked out following Section 2, is
$X=\operatorname{Spec}(\mathbb{C}[x,y,z]/(4xz+y^{2}+8z^{3},xy+4yz^{2},x^{2}+4y^{2}z+12xz^{2}+22z^{4}))\,$, (3.36)
@@
Based on the degree of $z$ being $2$, we can use the jet scheme $J_{10}(R)$ to determine the coefficients of the Schur limit of the Hilbert series of the arc space up to and including order $q^{12}$. Again utilizing Macaulay2, we find that the leading coefficients are
(1,0,1,2,4,5,9,12,21,29,44,60,91).(1,0,1,2,4,5,9,12,21,29,44,60,91)\,. (3.38)
@@ 4 Macdonald index
Effectively, this way of introducing an extra grading is tantamount to counting the number of letters needed to construct a (basis) vector in $\textrm{gr}V$ endowed with grading $w$. When there is more than one type of letter, then the relation may not respect this grading based on the degree of letters. For example, when we have $x^{2}=y^{3}$ with $w(x)=2$ and $w(y)=1$, the LHS has $w=4$ whereas RHS has $w=3$. The relation implies that we should not count both.
Our prescription is simply that whenever the relation mixes the $T$-grading, drop the states with the highest $T$-weight. This is realized by equation (4.3). For example, suppose the relation identifies the states with $T$-weight of $m$ and m1m-1. Then we see that such a vector of $T$-weight $m$ should be absent in $\textrm{gr}V^{(m)}=V_{m}/V_{m-1}$ from the relation. However such a vector is not removed in $\textrm{gr}V^{(m-1)}=V_{m-1}/V_{m-2}$ since the relation is in $V_{m}$ but not in $V_{m-1}$.
@@ 5.1 $\mathbb{C}^{2}/\mathbb{Z}_{2}$
We can see that once $X$ is reduced, we obtain the $\mathbb{C}^{2}/\mathbb{Z}_{2}$ singularity as a variety. We can use Macaulay2 to determine the Schur limit of the Hilbert series of the $7$th jet scheme, which should overlap with the Schur limit of the Hilbert series of the arc space up to and including the $q^{8}$ term. We find:
$\operatorname{HS}_{q,q}(J_{\infty}(R))=1+3q+9q^{2}+19q^{3}+42q^{4}+81q^{5}+155q^{6}+276q^{7}+486q^{8}+\cdots\,.$ (5.5)
@@
$\operatorname{PE}\left[\frac{q-q^{3}}{(1-q)(1-q^{3})}\chi_{3}\right]\,.$ (5.6)
We see that these match perfectly at least up to and including order $q^{8}$. While it would be useful to go to higher orders to provide a more convincing demonstration of our conjecture in equation (1.4), Macaulay2 has difficulty finding Gröbner bases for the higher jet schemes. In the recent work [5], the authors proved that the arc space of our affine scheme indeed gives the Schur and Macdonald indices.
In this analysis, we have chosen to introduce a degree-two coordinate $w$ corresponding to the stress-tensor of the 4d SQFT for convenience, however, the scheme $X$ is isomorphic to the scheme
$\operatorname{Spec}\left(\mathbb{C}[x,y,z]/((x^{2}+y^{2}+z^{2})x,(x^{2}+y^{2}+z^{2})y,(x^{2}+y^{2}+z^{2})z)\right)\,$, (5.7)
where the indeterminates have the same degrees as before. It is necessary (see [60] for more details) for our affine scheme to be bifiltered if it is to reproduce the Macdonald index. We can specify the bifiltration in the ring in equation (5.3) by inheriting it from the bigrading
$\operatorname{deg}(x)=\operatorname{deg}(y)=\operatorname{deg}(z)=(1,1)\,,\qquad\operatorname{deg}(w)=(2,1)\,$, (5.8)
on $\mathbb{C}[x,y,z,w]$. We can then push this bifiltration through the isomorphism to the ring in equation (5.7), however, this bifiltration cannot be inherited from a bigrading structure on the ring $\mathbb{C}[x,y,z]$. Therefore, we choose to include a coordinate $w$ corresponding to the stress-tensor explicitly to make the choice of bifiltration clear.
@@
$\operatorname{HS}_{q,q}(J_{\infty}(R))$ $=1+3q+9q^{2}+22q^{3}+51q^{4}+105q^{5}$ (5.10)
$\qquad\qquad+212q^{6}+402q^{7}+744q^{8}+1326q^{9}+2316q^{10}+\cdots\,$, (5.11)
@@
up to and including the $q^{10}$ term. Therefore, we conjecture that the following affine scheme captures the Macdonald and Schur indices of the $(A_{1},D_{2n+1})=\mathcal{D}_{2n+1}(SU(2))$ theory:
@@ 5.2 $\mathbb{C}^{2}/\mathbb{Z}_{n+1}$
which is known to have $\mathbb{C}^{2}/\mathbb{Z}_{3}$ as its Higgs branch. The associated vertex operator algebra for this theory is given by the Bershadsky–Polyakov (BP) W-algebra [21, 70], whose connection to Argyres–Douglas theories was found in [39]. The BP algebra has strong generators of dimensions $1,3/2,3/2,2$, and its Zhu’s algebra $R_{V}$ is identical to our $R$ so that the associated scheme $X_{V}$ is identical to our affine scheme XX. [^18] There is a conjectural form of the Zhu’s algebra written in [84] for the $(A_{1},A_{2n+1})$ theory, which is slightly different from what we have. This gives a different Hilbert series which does not match with the Schur index.